A gas occupies a volume of 400.0 ml at 29⁰C and 1.15 atm if the pressure is held constant by how much should the temperature be decreased so that the volume decreases to 361.6 ml


1
Expert's answer
2022-06-06T17:04:04-0400

Given:

V1 = 400.0 mL

T1 = 29°C + 273.15 = 302.15 K

P1 = P2 = 1.15 atm = constant

V2 = 361.6 mL

T2 = unknown


Formula: V1 / T1 = V2 / T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the resulting temperature, solve the equation for T2:

T2 = V2T1 / V1

T2 = (361.6 mL × 302.15 K) / (400.0 mL) = 273.14 K

T2 = 273.14 K (or 0°C)

ΔT = T1 − T2 = 29°C − 0°C = 29°C

Therefore, the temperature should be decreased by 29°C


Answer: The temperature should be decreased by 29°C